OpenStudy (anonymous):

solve this using substitution x=2y-2 3y=x+6

OpenStudy (a1234):

x=2y-2 3y=x+6 3y = 2y - 2 + 6 y = 4 x = 2(4) - 2 x = 8 - 2 x = 6 (6, 4) See if the answer works.